1. It is known that the average price for a cappuccino in a restaurant is R15.40 with a standard deviation of R1.20. What is the probability that the sampling error for the mean is less than 0.20 for a sample of 32 restaurants?
1
Expert's answer
2020-10-05T18:05:55-0400

CI =X̄ ± Z×s/"\sqrt n" =15.4 ±Z*1.2/"\sqrt{32}"

Z*1.2/"\sqrt{32}" <0.2

0.21Z<0.2

Z<1, so using table probability is near 70%